__6__ ___4____
(2x-3)= (9x-5)

Answer by checkley75(3666)

6/(2X-3)=4/(9X-5) NOW CROSS MULTIPLY & WE GET:
4(2X-3)=6(9X-5)
8X-12=54X-30
8X-54X=-30+12
-46X=-18
X=-18/-46
X=9/23 ANSWER.